OpenStudy (aaronandyson):

13 hrs 20 min in hours

OpenStudy (rational):

13 hrs 20 min is same as 13 hrs + 20 min

OpenStudy (rational):

Next use 1 min = 1/60 hour

OpenStudy (rational):

13 hrs + 20 min = 13 hrs + 20*1/60 hrs = 13 hrs + 1/3 hrs = 13 1/3 hrs

OpenStudy (aaronandyson):

? My book says 3/40? ?

OpenStudy (rational):

attach the screenshot if psble

OpenStudy (aaronandyson):

I don't go to a virtual school you see.

OpenStudy (rational):

if u have cam, you may take a picture and attach

OpenStudy (aaronandyson):

I don't have one either.